A barium hydroxide solution is prepared by dissolving 1.91 g1.91 g of Ba(OH)2Ba(OH)2 in water to make 25.2 mL25.2 mL of solution. What is the concentration of the solution in units of molarity?

concentration:

M

The barium hydroxide solution is used to titrate a perchloric acid solution of unknown concentration. Write a balanced chemical equation to represent the reaction between barium hydroxide and perchloric acid.

chemical equation:

If 23.2 mL23.2 mL of the barium hydroxide solution was needed to neutralize a 2.49 mL2.49 mL aliquot of the perchloric acid solution, what is the concentration of the acid?

concentration:
Answer #1

Ans :

Mol Ba(OH)2 = mass / molar mass

= 1.91 g / 171.34 g/mol = 0.0111 mol

Molarity = mol solute / volume solution(L)

putting values :

= 0.0111 mol / 0.0252 L

= 0.442 M

The chemical equation is given as :

Ba(OH)2 + 2HClO4 = Ba(ClO4)2 + 2H2O

Using the relation : M1V1 / n1 = M2V2 / n2

where M1 = 0.442 M

V1 = 23.2 mL

n1 = 1 mol

V2 = 2.49 mL

n2 = 2 mol

putting values :

0.442 M x 23.2 mL = (M2 x 2.49 mL) / 2

M2 = 8.24 M

So concentration of acid is 8.24 M.